APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday September 20, 2004 (MJD 53269) ---=== OBSERVING TEAM ===--- Steph Snedden (swing) Atsuko Nitta John Barentine The Amazing Mr. Loomis (support) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Imaging of 72 S from -174 < lambda < -126 Finished Plate 1916 Everything was hunky-dory for most of the night until network problems flummoxed us late (see Problems section). ---=== OBSERVING LOG ===--- Swing: ----- The absorbent cloths were moved to the corners of the lower enclosure, out of the way of the racks. The Grade 2 N2 bottle (for imager LN2 pressurization) is getting close to 500 PSI on the gauge, and will need to be changed before too long. There are no leaks and we've been running just above 500 PSI for the dark run, so there's no apparent rush. Spectro focus check done, and b1 was moved a small amount. Details at the end of the log. The rest of afternoon checkout was uneventful. Night: ------ 01:05Z Fans on, louvers open. Rolled the enclosure off. 02:00Z Atsuko noted that the M1 temperature sensors showed the primary was supercooling. Rather than take the chance, we pre-emptively closed. 03:05Z Opened again after the mirror temperatures stabilized. Force filled the spectrographs again because closing so soon after the last force fill resulted in a large amount of boil-off. 03:20Z Loaded Cart 1, Plate 1916. Stars on goToField after clearing offsets. Finished the plate in three exposures. 05:20Z Switched to imaging, stripe 72 S. Run 4843: Stripe 72 S, startLambda=-176. Seeing was good from the beginning. In focus by frame 17, skippy happy by frame 13. Ramp complete by Lambda = -173.8, Frame 19. Cloud camera happily reported photometric conditions throughout; seeing was good to excellent. Last science frame is 338 at lambda=-126. Run 4844: Stripe 72 N, startLambda=-166. By Frame 11, most of the chips went blue on the scrolling displays (all but dewar 1) -- maybe a nearby bright star? Evertyhing was fine again by Frame 12. Seeing, focus and skippy were good from the beginning. First science frame was Frame 31 at lambda=-162. Network problem 09:10-10:45Z. (see Problems section). We continued collecting imager data (but no PT) during this time. However, skippy did not run again after Frame 12; we finally got it back around Frame 133. Ended the drift at 10:42Z when clouds started moving in. We think this whole run should be declared bad because there were only about 30 useful frames before the network died (and we could no longer verify astrometry or photometricity). Run 4845: Bias run on the telescope. During this time, I lost my X session on commish altogether, and Watcher appears to have died (see Problems section). Gave up with < 1 hour until twilight and clouds. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4843 05:26Z 08:53Z 72 S -175.55 -123.45 362 science 4844 08:57Z 10:41Z 72 N -165.65 -139.57 187 science 4845 10:44Z 11:30Z -- -- -- 44 On-telescope bias (times are only approximate) ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 4843 13 1 25 -6.100 0.209 0.600 0.085 0.00658 96 59 4843 22 4 82 -6.600 0.604 0.600 0.531 0.00621 96 59 4843 75 4 65 -7.800 0.443 0.400 0.454 0.00385 97 58 4843 80 4 68 -7.400 0.178 0.500 0.121 0.00441 97 58 4843 128 4 76 -6.300 0.392 0.500 0.390 0.00479 98 57 4843 181 4 74 -6.800 0.244 0.300 0.175 0.00491 100 56 4843 234 4 69 -8.200 0.170 0.200 0.101 0.00328 103 55 4843 285 4 58 -7.500 0.154 0.100 0.099 0.00258 107 54 4844 12 1 20 -6.000 0.102 0.100 0.081 0.03347 222 79 4844 133 4 62 -5.800 0.259 -2.000 0.458 0.02624 209 76 4844 133 4 62 -5.800 0.259 -2.000 0.458 0.02624 209 76 4844 157 4 46 -4.500 0.122 -2.500 0.089 0.02371 207 76 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 4843 34 1 59 84 25 -1.988 0.154 -0.674 0.101 0.00227 4843 34 1 59 84 29 -1.567 0.191 -0.564 0.171 0.00189 4843 34 1 59 84 31 -1.108 0.203 -0.830 0.195 0.00279 4843 34 1 59 84 27 -0.583 0.267 -0.850 0.165 0.00286 4843 34 1 59 84 32 -0.309 0.217 -0.827 0.160 0.00278 4843 51 1 59 83 15 -1.041 0.136 -1.033 0.125 0.00347 4843 51 1 59 83 30 -0.763 0.143 -0.986 0.135 0.00331 4843 51 1 59 83 26 -0.256 0.138 -1.054 0.164 0.00354 4843 51 1 59 83 26 0.097 0.209 -1.034 0.167 0.00348 4843 51 1 59 83 33 0.563 0.171 -1.094 0.142 0.00368 4843 85 1 58 83 25 -1.453 0.154 -0.604 0.217 0.00203 4843 85 1 58 83 21 -1.173 0.162 -0.501 0.128 0.00169 4843 85 1 58 83 27 -0.727 0.156 -0.492 0.109 0.00165 4843 85 1 58 83 30 -0.287 0.108 -0.599 0.190 0.00201 4843 85 1 58 83 20 0.117 0.137 -0.613 0.219 0.00206 4843 103 1 58 82 30 -1.604 0.296 -0.880 0.134 0.00296 4843 103 1 58 82 23 -1.058 0.193 -0.953 0.119 0.00321 4843 103 1 58 82 21 -0.566 0.183 -0.970 0.113 0.00326 4843 103 1 58 82 23 -0.245 0.217 -0.934 0.068 0.00314 4843 103 1 58 82 25 0.203 0.173 -1.029 0.122 0.00346 4843 140 1 57 81 19 -1.327 0.142 -1.193 0.094 0.00401 4843 140 1 57 81 28 -0.998 0.083 -1.145 0.127 0.00385 4843 140 1 57 81 23 -0.546 0.147 -1.220 0.134 0.00410 4843 140 1 57 81 28 -0.081 0.194 -1.156 0.116 0.00389 4843 140 1 57 81 30 0.259 0.165 -1.259 0.195 0.00423 4843 156 1 57 80 23 -1.229 0.143 -0.831 0.157 0.00280 4843 156 1 57 80 35 -0.833 0.170 -0.716 0.118 0.00241 4843 156 1 57 80 21 -0.431 0.184 -0.853 0.156 0.00287 4843 156 1 57 80 20 -0.042 0.181 -0.770 0.105 0.00259 4843 156 1 57 80 23 0.177 0.205 -0.634 0.250 0.00213 4843 191 1 56 79 29 -1.640 0.199 -0.504 0.101 0.00169 4843 191 1 56 79 21 -1.282 0.204 -0.451 0.097 0.00152 4843 191 1 56 79 24 -0.789 0.194 -0.514 0.204 0.00173 4843 191 1 56 79 30 -0.438 0.215 -0.574 0.210 0.00193 4843 191 1 56 79 22 -0.032 0.186 -0.597 0.188 0.00201 4843 210 1 55 77 24 -1.055 0.132 -0.493 0.137 0.00166 4843 210 1 55 77 21 -0.636 0.153 -0.420 0.094 0.00141 4843 210 1 55 77 27 -0.123 0.155 -0.521 0.128 0.00175 4843 210 1 55 77 16 0.263 0.193 -0.544 0.116 0.00183 4843 210 1 55 77 18 0.581 0.157 -0.514 0.112 0.00173 4843 244 1 55 75 22 -2.068 0.121 -0.295 0.139 0.00099 4843 244 1 55 75 24 -1.590 0.193 -0.327 0.104 0.00110 4843 244 1 55 75 20 -1.036 0.211 -0.425 0.195 0.00143 4843 244 1 55 75 27 -0.405 0.197 -0.560 0.153 0.00188 4843 244 1 55 75 25 -0.153 0.190 -0.507 0.178 0.00170 4843 261 1 54 74 29 -1.978 0.224 0.106 0.108 -0.00036 4843 261 1 54 74 30 -1.438 0.204 0.087 0.071 -0.00029 4843 261 1 54 74 27 -0.937 0.151 0.075 0.124 -0.00025 4843 261 1 54 74 22 -0.571 0.205 0.031 0.148 -0.00011 4843 261 1 54 74 27 -0.177 0.198 -0.051 0.162 0.00017 4843 297 1 53 72 31 -1.333 0.100 -0.157 0.148 0.00053 4843 297 1 53 72 28 -1.078 0.161 -0.062 0.145 0.00021 4843 297 1 53 72 23 -0.690 0.145 -0.050 0.149 0.00017 4843 297 1 53 72 23 -0.226 0.143 0.018 0.153 -0.00006 4843 297 1 53 72 27 0.284 0.190 -0.162 0.163 0.00055 4843 315 1 53 71 19 -0.593 0.139 -0.222 0.121 0.00075 4843 315 1 53 71 17 -0.435 0.123 -0.054 0.109 0.00018 4843 315 1 53 71 29 0.133 0.139 -0.152 0.183 0.00051 4843 315 1 53 71 27 0.526 0.169 -0.176 0.206 0.00059 4843 315 1 53 71 29 0.913 0.181 -0.257 0.170 0.00086 4844 74 1 78 326 24 -1.438 0.258 -0.568 0.120 0.00191 4844 74 1 78 326 25 -0.877 0.192 -0.547 0.159 0.00184 4844 74 1 78 326 21 -0.371 0.294 -0.573 0.213 0.00193 4844 74 1 78 326 28 -0.079 0.242 -0.644 0.112 0.00216 4844 74 1 78 326 32 0.278 0.258 -0.671 0.104 0.00226 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 28619, plate 1559 ------- 23:34 28619 10.0 flat 23:36 28620 2.0 arc 23:40 28621 2.1 arc 23:42 28622 2.1 arc 00:13 28623 2.1 arc 00:16 28624 2.1 arc 00:22 28625 2.1 arc 00:24 28626 2.1 arc ----- sequence 28627, plate -9999 ------- 00:46 28627 0.0 bias ----- sequence 28628, plate -9999 ------- 02:51 28628 0.0 bias ----- sequence 28628, plate 1916 ------- 03:36 28629 10.0 flat 03:38 28630 2.0 arc 04:01 28631 900.1 target 4.6 5.2 7.2 12.4 04:12 28632 1200.1 target 7.3 6.2 10.5 15.6 04:46 28633 1200.1 target 7.8 6.9 9.1 13.2 04:50 28634 10.0 flat 04:52 28635 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:45Z 1 1916 0.97 0.0016 51.21 0.0016 180.69 -0.0118 1.000270 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL7794-7800 Stays: JL7801-07 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:44Z 1 1916 1.00027 -2665 -980 50 1 51.2 14.3 8 233 gdr 1.5 06:12Z imager 1.00000 0 -319 -211 83 58.1 14.3 10 263 r' 1.0 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:16Z 57 0 14 249 (WSW) 280 - Clear 23:48Z 57 0 11 241 (WSW) 328 - Clear 00:19Z 57 0 12 239 (WSW) 370 - Clear 00:51Z 57 0 9 267 (W) 480 - Clear 01:46Z 57 0 11 239 (WSW) 584 - Clear 02:16Z 57 0 11 236 (SW) 581 - Clear 02:47Z 57 0 11 240 (WSW) 563 - Clear 03:17Z 57 0 14 238 (WSW) 617 - Clear 03:51Z 57 0 7 224 (SW) 611 - Clear 04:24Z 57 0 12 241 (WSW) 512 - Clear 04:56Z 57 0 11 242 (WSW) 554 - Clear 05:29Z 57 0 8 262 (W) 605 - Clear 06:01Z 57 0 5 252 (WSW) 499 - Clear 06:33Z 57 0 12 264 (W) 549 - Clear 07:05Z 57 0 11 251 (WSW) 493 - Clear 07:37Z 57 0 14 240 (WSW) 530 - Clear 08:10Z 57 0 16 222 (SW) 504 - Clear 08:42Z 57 0 15 207 (SSW) 520 - Clear 09:12Z 57 0 19 223 (SW) 532 - Clear 09:42Z 57 0 14 218 (SW) 532 - Clear 10:12Z 57 0 18 216 (SW) 532 - Cloudy 10:43Z 57 0 18 216 (SW) 532 - Cloudy 11:30Z 57 0 18 219 (SW) 532 - Cloudy 12:02Z 57 0 20 214 (SW) 532 - Cloudy 12:38Z 57 0 12 210 (SSW) 532 - ---=== TELESCOPE STATUS ===--- 01:05Z Fans on, louvers open, enclosure off 02:00Z Enclosure on (high humidity) 03:05Z Enclosure off 11:30Z Enclosure on, fans off, louvers closed Telescope stowed at 121,30 with Cart 4, Plate 1917 mounted Counterweights at 250, spectro autofill reconnected No interlocks bypassed Dewars at 12:00Z Spectro 108lbs 25psi Imager 132lbs 21psi ---=== SOFTWARE USED ===--- IOP/SOP: v3_140_0 Watcher: v2_29_0 MCP: v5_25_0 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v1_93 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5330 0.5290 Axial B 0.6100 0.6100 Axial C 0.9240 0.9250 Trans D -8.1760 -8.1740 Lateral E 10.0457 10.0457 Lateral F 11.5951 11.5951 GALILS Commanded: 4950. -2250. 2550. -13750. -7850. -7900. Actual: 4949. -2231. 2533. -13737. -7862. -7903. SETMIR VALUES PriDesOrient: 0.00 -5.30 1.00 57.00 -288.80 PriOrient: 0.00 -5.32 1.32 56.82 -289.65 SECONDARY: ---------- Focus: 0.00 Air Temp.: 14.3 Alt.: 30.072645 MIGS TONIGHT NOMINAL Axial A 1.5190 1.5190 Axial B 1.0960 1.0970 Axial C 1.1100 1.1110 Trans D 0.0030 0.0020 GALILS Commanded: 1635649. 1562729. 1552840. 2850. -10650. Actual: 1635720. 1562731. 1552929. 2828. -10665. SETMIR VALUES SecDesOrient: 1257.00 11.62 -25.00 100.00 83.30 SecOrient: 1256.96 11.64 -24.98 100.09 82.85 ---=== PROBLEMS IN DETAIL ===--- Spectro focus check: ------------------- T=12.9C, Cartridge 2 Before adjust sp1 Coll_motor_A -3698 Coll_motor_B -2727 Coll_motor_C -4799 sp2 Coll_motor_A -1976 Coll_motor_B -3194 Coll_motor_C -3974 B1 was moved -11 deg. This did not change the coll motor numbers, of course. Intranet Outage: ---------------- Just after 09:10Z we appeared to have lost a number of internal network connections around the site. MOP crashed on sdssmth and required killing windows to exit the software. hoggpt quit updating cloud camera images. We could not ping consistently between computers on site, but could ping domains outside apo.nmsu.edu. We could not re-initialize the connection to the PT from sdssmth. Throughout this time, we continued imaging (i.e., gangs were bring transferred properly). We got Craig out of bed at 05:30am EDT via cell phone, who graciously talked us through a determination of the problem. As he put it, "NMSU stopped advertising the apo.nmsu.edu delegation." While our local network was, in fact, fine, the "rest of the world" didn't know that apo.nmsu.edu existed during this time. This MAY have affected incoming mail and even mail between machines on site (because of the required lookup of apo.nmsu.edu addresses); as of this writing (12:10Z), despite having sent out the PT log, plugging email, etc., none of us have received those messages in our galileo accounts. We will leave a paper note for the plate pluggers in case our email vaporizes in the meantime. Last but not least, we cannot ssh to tapehost@omaha ("ssh: connect to host omaha port 22: No route to host"). What Is It With These Computers, Anyway?: ----------------------------------------- As we ended the on-telescope bias run near 10:45Z, my (JCB) X session on commish from the terminal in the control room died. I lost control of all the windows. Right around the same time, Atsuko reported that Watcher had died on her landru session; it was also gone on monx3. SN group tape drive: -------------------- FYI, in the cluster of tape drives in the computer room beloning to the SN group, the drive called "col3" has a cleaning light blinking. We ran a cleaning tape in that drive. TPM Errors: ----------- dewar (Imager Dewar Scale Communications Fault:: 2004-09-21 11:28:21Z) m2 mig (Communications Fault: 2004-09-21 11:29:21Z)